OpenStudy (anonymous):

(x^3) to the 3/4 power. simplify. . Express the product as a radical expression.

OpenStudy (anonymous):

\[x^{3^{3/4}}=?\]Is that you want

OpenStudy (anonymous):

?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but the x^3 is in parentheses

OpenStudy (anonymous):

i believe its \[x^3 \sqrt[4]{x^3}\]

OpenStudy (anonymous):

not sure though

OpenStudy (anonymous):

ok then using the property\[(a^{n})^{m}=a^{mn}\] therefore,\[x^{3^{3/4}}= x^{2*3/4}= x^{3/2}=\sqrt[2]{x^{3}}\]

OpenStudy (anonymous):

This does not simplify as an expression... \[\huge (b^m)^n=b^{m\cdot n}\] \[(b^3)^3=b^3*b^3*b^3=(b*b*b)*(b*b*b)*(b*b*b)=\large b^{3\ \cdot\ 3}=b^9\]

OpenStudy (anonymous):

shubham.bagrecha thats not it im afraid. the answer choices are: \[x^3[\sqrt[4]{x^3}\] \[x^2\sqrt[4]{x^3}\] \[x^2\sqrt[4]{x}\] \[x\sqrt{x}\]

OpenStudy (anonymous):

You can rewrite it but that's technically not "simplified" :-)

OpenStudy (anonymous):

I take beef with your question's author :P

OpenStudy (anonymous):

okay so how would i rewrite it so it looks like the answer choices i have posted above?

OpenStudy (anonymous):

Ah! I see your answer

OpenStudy (anonymous):

what is x^(1/2) ;-)

OpenStudy (anonymous):

Same as \(\sqrt{x}\), no?

OpenStudy (anonymous):

no its just x raised to the power of 1 divided by two

OpenStudy (anonymous):

\(\frac{3}{2}\) is the same as 1\(\frac{1}{2}\)

OpenStudy (anonymous):

And dividing by a power is the same as taking the n-th root, by definition of what a root is.

OpenStudy (anonymous):

okay so whats \[x^3\] raised to the power of 3 divided by 4

OpenStudy (anonymous):

. Express the product as a radical expression.

OpenStudy (anonymous):

As @shubham.bagrecha said, x\(^{3/2}\) Write it out expanded like did initially! It'll help ;-)

OpenStudy (anonymous):

Does writing it like this help? \[\sqrt{x*x*x}\]

OpenStudy (anonymous):

What is \(\sqrt{x*x}\) = ? ;-)

OpenStudy (anonymous):

x right?

OpenStudy (anonymous):

Or with 2's \(\sqrt{2*2*2}\)

OpenStudy (anonymous):

we can write it as \[x^{2/2}*x^{1/2}=x^{3/2}\] =\[=x*\sqrt{x}=x^{3/2}\]

OpenStudy (anonymous):

Yep that's right

OpenStudy (anonymous):

So the fourth option is the asn.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Again, it's not technically "simplified" just rewritten. Simplification just means having everything in base form, in layman's terms. But good enough. This is trying to see if you understand the concept of what to do with fractional exponents. Rules of exponents.